OpenStudy (anonymous):

f(x) = square root of quantity x minus two. and g(x) = square root of quantity x plus six. I'm pretty sure it's 4 units but how do I figure out if it's to the left or right?

OpenStudy (anonymous):

These are the answers: Shift the graph of f left 8 units. Shift the graph of f right 4 units. Shift the graph of f left 4 units. Shift the graph of f right 8 units.

OpenStudy (anonymous):

These are the directions: Describe how to transform the graph of f into the graph of g.

OpenStudy (anonymous):

@mathmath333 any help?

OpenStudy (anonymous):

@animal_lover36 @campbell_st @nnsfnjsnfj @TheSmartOne @woooooooooowwwww @zuleimasainz HELP!!??

OpenStudy (mathmath333):

\(\large\tt \color{black}{f(x)=\sqrt{x-2},g(x)=\sqrt{x+6}}\) \(\large \tt \color{black}{\text{are these two ?}}\)

OpenStudy (anonymous):

Yes @mathmath333 that's right

OpenStudy (anonymous):

I got help from someone else and the answer we got was 8 units to the left is that correct? @mathmath333

OpenStudy (mathmath333):

Shift the graph of f left 8 units.

OpenStudy (anonymous):

Thank you :)